Find the volume of the solid generated by revolving the region bounded by the curves x=y^2, x=2, y=0; about y=0


1
Expert's answer
2021-05-28T08:24:48-0400

We have given the curve,


"x = y^2,x = 2,y = 0"


Volume of solid can be calculated as,


"V = \pi \int_{0}^{2}(\sqrt{x})^2dx"

"V = \pi \int_{0}^{2}xdx"

"V = \pi \times\dfrac{4}{2}"

"V = 2 \pi"